### v1

# Improved treatment of relativistic effects in linear augmented plane wave

Abstract
We examine the influence of the relativistic effects within the linear augmented plane wave method (LAPW) with the potential of general shape for solids and suggest a few ways to account them more accurately: (1) we introduce new radial wave functions based on two actual radial solutions of the Dirac equation for j=l1/2j=l-1/2 and $j=l+1/2$ one-electron states; (2) the canonical LAPW matrix elements for the spherically symmetric component of the potential, assuming non-relativistic radial wave functions, should be corrected; (3) we argue that for a realistic spin-orbit energy splitting of the semicore $6p$-states the spin-orbit interaction constant $\zeta(p)$ should be calculated with the $6p_{3/2}$ radial component; (4) in cases when two $j=l\pm 1/2$ components are occupied (for example for the $6p$ states of actinides) the electron density, associated with the small components of valence electrons, can be taken into the calculation scheme. We demonstrate that the new treatment for the relativistic effects is capable to change the equilibrium lattice constant up to 0.15 Å and the bulk modulus up to 26 GPa. We find that the electron density of valence electrons at the nucleus increases by 2.34.32.3-4.3 times due to the inclusion of small components, which can be essential for precise description of the potential and density close to the nuclear region, important for nuclear spectroscopies. In contrast to the common believe that in plain band structure treatment UO2 is a metal, we show that in the presence of the spin-orbit coupling UO2 has a small gap of forbidden states (0.20.40.2-0.4 eV) at the Fermi level, where the highest occupied and the lowest unoccupied $5f$ bands slightly overlap, as in calculations of the conduction and valence band in solid Ge.
I Introduction
Nowadays band structure calculations become a powerful tool of investigation of complex materials proving its efficiency for many solids and capable of predicting their properties. The accuracy of such calculations increases with every year and there is a constant demand for even better precision and performance. As shown in benchmark calculations [1], various band structure methods generally result in the same or close final results (for example, the equilibrium lattice constants, bulk moduli etc.). Neverthelss, there is a class of materials where the description of solids is less certain and encounters difficulties. Such materials include, in particular, heavy elements – actinides, situated in the end of the Periodic Table, which heave eighty or more core electrons, experiencing relativistic effects. In the literature there are several successful studies of band structure of actinides. First works were based on the linear augmented muffin-tin method (LMTO) [2, 3, 4, 5], followed by studies [6, 7] carried out with the full potential linear augmented plane wave method (FLAPW) [8, 9, 10, 11], considered as one of the most precise band structure methods. Additional complexity of the subject is related to the fact that the electrons, belonging to the incompletely filled $5f$ shell of actinides exhibit competition between itineracy and localization [12, 13]. This competition can lead to nontrivial magnetic and other correlation effects [14, 15, 16, 17, 18].
However, even in the one-electron case the standard band treatment faces difficulties. With the nuclear charge $Z>90$, and the total number of electrons close to one hundred, the density functional theory (DFT) in the nuclear region is close to the regime of extreme electron densities [19, 20], absent for other materials. In addition, the relativistic effects of these elements are maximal. For example, the small component $Q$ of the Dirac equation, which can be safely ignored almost for all elements, is very large there. A proper description of electron structure also requires a consideration of the finite-size nucleus, which otherwise leads to singularities at the origin [21].
In the canonical FLAPW approach one uses so called scalar relativistic approach based on the works of Koelling and Harmon in [22], and MacDonald, Pickett and Koelling in Ref.[23]. (We consider it in detail below in Sec. II.1.) In a further development, it was proposed that the FLAPW basis set be enriched with local atomic functions [24, 25], which can be fully relativistic, i.e. taken from the solution to the radial Dirac equation. In Ref. [24] the relativistic $p_{1/2}$ local orbitals were added in the second variation step of the FLAPW calculation of elemental thorium, which turned out to significantly improve the stability and precision of band calculations. Two variational procedures are often used as a time-saving computational scheme in the FLAPW method when the spin-orbit (SO) coupling is included. In the first step only the scalar relativistic part of the Hamiltonian is diagonalized, whereas in the second variation step the SO coupling matrix is constructed and then diagonalized in a smaller basis set, consisting of a limited number of low lying eigenfunctions obtained on the first step and additional local atomic orbitals [24, 25]. In Ref. [24] only the $p_{1/2}$ relativistic atomic functions were used, while in Ref. [25] the method was extended to include other local relativistic functions and their combinations.
The proposed corrections [24, 25] through the relativistic local orbitals in the second variation step being highly effective in practice, are based on the idea to increase the convergence and effectiveness of the basis set. It does not improve the relativistic characteristics of canonical LAPW band functions. As shown later in Sec. II the present scalar relativistic LAPW method [22, 23], when applied to heavy elements like actinides and transactinides, demonstrates deviations of the averaged relativistic radial functions from the canonical scalar relativistic ones (Sec. II.1) and require small corrections in expressions for matrix elements (Sec. II.2). In addition, the second variation step performed on a small number of secondary basis functions (which is usually the case) can be considered as a perturbative treatment of the spin-orbit coupling [6].
The aim of the present study is to increase the accuracy of the relativistic effects within the full electron full potential LAPW method (FLAPW) method for heavy elements as much as possible, while keeping the general scheme of the method unchanged. For that purpose (1) we introduce new basis functions, obtained from two independent solutions of the Dirac equation; (2) we reconsider the matrix elements of the method, explicitly avoiding the use of hidden non-relativistic relations; (3) we correct the calculation of the spin-orbit (SO) coupling constant for $6p$ semicore states, based on the comparison the energy splittings between $6p_{1/2}$ and $6p_{3/2}$ components. In the present study we do not apply the second variation procedure for the SO-coupling. We use the direct treatment of the SO coupling in the full LAPW basis set thereby avoiding the approximations associated with the second variation step [10]. In this respect it is worth noting that in some relativistic calculations the authors tend to avoid the inclusion of the SO interaction [19, 20, 5], which, as discussed in Sec. II.3 below, can be connected with the overestimated values of SO coupling [6]. Finally, (4) we include the small components $Q$ in the calculations, which play an important role in the neighborhood of the nuclear region. These effects are considered and discussed in Sec. II.1, Sec. II.2, Sec. II.3, and Sec. II.4, correspondingly. In Sec. III we briefly review the results of our calculations for Ac, Th, ThO2 and UO2, and finally in Sec. IVwe discuss main conclusions and findings of our work. In our study we use various variants of the DFT functionals (see Sec. III below), which allows us to test the accuracy of the calculations.
II Method
In the LAPW method [8, 9, 10], widely used for studies of bulk materials, the space is partitioned in the region inside the nonoverlapping muffin-tin (MT) spheres and the interstitial region (IR). The basis functions $\phi_{j}(\vec{k},\,\vec{R})$, where $j=1,2,...,N_{b}$, are given by
$\phi_{j}(\vec{k},\,\vec{R})=\left\{\begin{array}[]{ll}v^{-1/2}\,exp(i(\vec{k}+\vec{K}_{j})\vec{R}),&\vec{R}\in IR\\ \sum_{l,m}{\cal R}_{l,m}^{j,\alpha}(r,\,E_{l})\,Y_{l,m}(\hat{r}),&\vec{R}\in MT(\alpha)\end{array}\right$. (1)
where $\vec{K}_{j}$ refers to the reciprocal lattice vector $j$, $v$ is the unit cell volume, $Y_{l,m}$ are spherical harmonics [26] and the radial part is given by
${\cal R}_{l,m}^{j,\alpha}(r,\,E_{l})=A^{j,\alpha}_{l,m}\,u_{l}(r,E_{l})+B^{j,\alpha}_{l,m}\,\dot{u}_{l}(r,E_{l}).\quad$ (2)
Here the index $\alpha$ refers to the type of atom (or MT-sphere) in the unit cell, the radius $r$ is counted from the center $\vec{R}_{\alpha}$ of the sphere $\alpha$ (i.e. $\vec{r}=\vec{R}-\vec{R}_{\alpha}$). Radial functions $u_{l}(r,E_{l})$ are solutions in the spherically averaged crystal potential computed at the linearization energy $E_{l}$, and $\dot{u}_{l}(r,E_{l})$ is the derivative of $u_{l}$ with respect to $E$ at $E_{l}$. The coefficients $A^{j,\alpha}_{l,m}$ and $B^{j,\alpha}_{l,m}$ are found from the condition that the basis function $\phi_{j}$ is continuous with continuous derivative at the sphere boundary, i.e. at $r=R_{MT}^{\alpha}$ ($R_{MT}^{\alpha}$ is the radius of the MT-sphere $\alpha$). The coefficients $A^{j,\alpha}_{l,m}$ and $B^{j,\alpha}_{l,m}$ in Eq. (2) are related to the standard LAPW quantities $a^{j}_{l}$, $b^{j}_{l}$, expressed only through the spherical Bessel functions $j_{l}$ and the radial solution $u_{l}$ (and its derivatives) at $r=R_{MT}^{\alpha}$.
II.1 Explicitly averaged radial basis wave functions
Initially, the functions $u_{l}(r,E_{l})$ in Eq. (2) were considered as the solutions of the Schrödinger equation in the spherically symmetric ($L=0$) component of the total potential. Later, it appeared that some relativistic effects can be included in the so called scalar relativistic approach [22, 23]. Below we discuss the canonical radial functions introduced by Koelling and Harmon in [22], later justified by the procedure described by MacDonald, Pickett and Koelling in Ref.[23], and compare them with new radial functions that are more closely related to the Dirac solutions.
The standard LAPW radial basis functions thus are given by an average
$P^{av}_{l}(r)=\frac{l}{2l+1}P_{l}(r)+\frac{l+1}{2l+1}P_{-l-1}(r),$ (3)
where $r$ is radius and $P_{l}$, $P_{-l-1}$ are the large ($L$) components of the Dirac solutions $P_{\kappa^{L}}$ for $\kappa^{L}=l$ (j=l1/2j=l-1/2), and $\kappa^{L}=-l-1$ ($j=l+1/2$), correspondingly. (Here $\kappa^{L}$ stands for the index $\kappa$ of 2-spinors for the large component [21], Appendix A.) However, in practice the large components $P_{l}$ and $P_{-l-1}$ are not calculated. In Ref. [23] assuming that
$\frac{d}{dr}(\delta P(r))=\frac{d}{dr}\left(P_{-l-1}(r)-P_{l}(r)\right)=0,$ (4)
an effective system for two coupled differential equations was derived. Then the LAPW radial basis function is
$P^{KH}_{l}(r)=\left.P^{av}_{l}(r)\right|_{\delta P^{\prime}(r)=0},$ (5)
i.e. the function (3), provided that the condition (4) is fulfilled. The second auxiliary function is an averaged small component $Q^{KH}_{l}(r)$ [23], given by Eq. (3b) of [23], i.e.
$Q^{KH}_{l}(r)=\frac{l}{2l+1}Q_{-l}(r)+\frac{l+1}{2l+1}Q_{l+1}(r),$ (6)
but except in the system of differential equations, it is not used. We recall that for the index $\kappa^{S}$ for the small ($S$) component $Q_{\kappa^{S}}$ we have $\kappa^{S}=-\kappa^{L}$ [21], Appendix A. As a result, $P^{KH}_{l}$, $Q^{KH}_{l}$ depend only on $l$ and include some relativistic effects. Although this approach has proved being efficient and practical, it has serious drawbacks when applied to heavy elements. In particular, Eq. (4) is only an approximation, and the averaging for $Q^{KH}_{l}$ in Eq. (5) is a formal procedure, for angular two-spinors $\xi_{-\kappa,m}$, associated with the Dirac small components $Q_{-l}$ and $Q_{l+1}$, have different angular dependencies [21], Appendix A. This is especially true for the $6p_{1/2}$ and $6p_{3/2}$ semicore states of all actinides. The problem is a very different radial behavior of $p_{1/2}$ and $p_{3/2}$ at small $r$, and below we consider it in more detail. It is well known that for the point nuclear case the Dirac solutions for $|\kappa|=1$ – that is, for $s_{1/2}$ and $p_{1/2}$ – have a singularity at the origin, which is absent for the $p_{3/2}$ radial functions. For the finite nuclear case, which is realized in our LAPW version, the singularity disappears ($P/r\sim r^{l}=r$) but the noticeable difference remains.
Figure 1: Radial Dirac functions $P_{j=1/2}(r)$ and $P_{j=3/2}(r)$ of the $6p_{1/2}$ and $6p_{3/2}$ semicore states of thorium, (a) close to the nuclear region, and (b) inside the MT-sphere.
To make the consideration on the radial part more concrete in the following we consider the case of the face centered cubic (fcc) lattice of elemental thorium (with the PBE exchange correlation functional [27]). In Fig. 1 we plot the radial dependencies $P_{j=1/2}(r)$ and $P_{j=3/2}(r)$ for the large components of $6p_{1/2}$ and $6p_{3/2}$ states at the same energy obtained by solving the Dirac equation in the self-consistent spherical potential. One clearly sees the different shape of large components both at the large radii close to the MT-radius and at the neighborhood of the nuclear region. Then we calculate the average radial function $P^{av}_{\ell=1}(r)$ for the $6p$-states using Eq. (3) explicitly. This numerically averaged function $P^{av}_{\ell=1}(r)$, as well as the conventional LAPW function $P_{l=1}^{KH}$, obtained by solving the KH differential equations [22, 23], are reproduced in Fig. 2. We see that the numerically averaged radial function $P^{av}_{\ell=1}$ remains different from $P_{l=1}^{KH}$ both at small and large radii, reflecting the approximate character of Eq. (4) and Eq. (5). In the following instead of $P_{l=1}^{KH}(r)$, and the corresponding energy derivative radial function $\dot{P}^{av}_{\ell=1}(r)=\partial P^{av}_{\ell=1}(r)/\partial E$, required by the LAPW method, we suggest to use the explicitly averaged functions $P^{av}_{\ell=1}$, $\dot{P}^{av}_{\ell=1}$ as the $6p$ LAPW basis set.
Figure 2: Radial basis function $P^{av}_{\ell=1}(r)$ (avD) and the canonical KH radial basis function $P^{KH}_{l=1}(r)$ [22, 23] of the semicore $6p$ states of thorium, (a) close to the nuclear region, and (b) inside the MT-sphere.
Similarly to the explicit procedure of averaging $p_{1/2}$ and $p_{3/2}$, described above, we can introduce new radial functions for $d$- and $f$- (and high $\ell$) states by using Eq. (3) for the independently calculated j=l1/2j=l-1/2 and $j=l+1/2$ Dirac radial functions. In the following we will refer to these directly averaged radial functions $P^{av}_{\ell}$ (large components) of the Dirac solutions as new basis functions (avD basis set) comparing their performance with the standard KH-basis functions $P^{KH}_{l}$. To characterize quantitatively the difference between $P^{av}_{\ell}$ and $P_{l}^{KH}$ ($\dot{P}^{av}_{\ell}$ and $\dot{P}_{l}^{KH}$) for $\ell>0$ we introduce the deviation quantities $\triangle P_{l}$ and $\triangle\dot{P}_{l}$, defined as
$\triangle P_{l}$ $=$ $\sqrt{\int_{0}^{R_{MT}}(P^{av}_{\ell}(r)-P_{l}^{KH}(r))^{2}},$ (7a)
$\triangle\dot{P}_{l}$ $=$ $\sqrt{\int_{0}^{R_{MT}}(\dot{P}^{av}_{\ell}(r)-\dot{P}_{l}^{KH}(r))^{2}},$ (7b)
Calculated $\triangle P_{l}$ and $\triangle\dot{P}_{l}$ for various elements and compounds are listed in Table 1. Note that the large component of $s_{1/2}$ functions in both treatments coincides, i.e. $P^{av}_{\ell=0}=P_{l=0}^{KH}$, $\dot{P}^{av}_{\ell=0}=\dot{P}_{l=0}^{KH}$.
Inspection of Table 1 shows that the largest difference between two functions (more than 10% of its norm) is found for $6p$-basis states of actinides. The differences between the $d$- and $f$- functions are of the order of only $10^{-3}$, but one should have in mind that Eq. (7a) and Eq. (7b) are integral. As we will see in Sec. II.3 even this small difference matters for the calculation of SO coupling constants $\zeta$, because in atomic units
$\zeta=\frac{1}{2c^{2}}\left\langle\frac{1}{r}\frac{dV}{dr}\right\rangle_{av}\sim\left\langle\frac{Z}{r^{3}}\right\rangle_{av},$ (8)
where $V(r)$ is the Coulomb potential. Therefore in this case the neighborhood of the nucleus, where the differences are visible, Fig. 3, contributes with considerably larger weight than the other regions. In general, however, the explicitly averaged radial functions (avD) for $6d$ ($P^{av}_{\ell=2}$), $5f$ ($P^{av}_{\ell=3}$) and higher $\ell$-states demonstrate a much more close correspondence with the KH-radial functions $P_{l=2}^{KH}$, $P_{l=3}^{KH}$, etc., because of weak presence of these functions in the nuclear region. We will return to this problem in Sec. II.3 below. Differences between the avD and KH radial basis functions for light elements such as oxygen in ThO2 or UO2 are negligible, Table 1.
Table 1: Deviations $\triangle P_{l}$ and $\triangle\dot{P}_{l}$, Eq. (7a) and Eq. (7b), between two radial basis functions: the canonical KH $P_{l}^{KH}$ ($\dot{P}_{l}^{KH}$) and the explicitly averaged Dirac (avD) functions $P_{\ell}^{av}$ ($\dot{P}_{\ell}^{av}$), Eq. (3), in fcc Th, fcc Ac, bcc Np, cubic ThO2 [Th(2), O(2)] and UO2 [O(3)] for $l>0$; $s$-functions coincide ($\triangle P_{l=0}<10^{-10}$).
$p$ $d$ $f$ $\ell>3$
[…]
II.2 Correction of LAPW matrix elements
The use of relativistic basis functions requires a modification of some matrix elements which are valid only in the non-relativistic limit. In particular, the matrix elements of the $L=0$ component of the potental, as written in Eq. (16a) and Eq. (16b) of Ref. [8] are exact only in nonrelativistic limit. This holds because in deriving the expressions for the matrix elements, the equality
$R_{MT}^{2}(\dot{u}_{l}(R_{MT})\,u^{\prime}_{l}(R_{MT})-\dot{u}^{\prime}_{l}(R_{MT})\,u_{l}(R_{MT}))=1,\quad$ (9)
is used. (Here, as before, the energy derivative $\dot{u}(r)=\partial u(r)/\partial E$ is defined in Rydberg energy units.) Eq. (9), explicitly quoted in Ref. [8] as Eq. (4), is exact only for the radial component $u_{l}$ of the Schrödinger equation in the spherically symmetric potential.
In general, the expression on the left hand side of Eq. (9) deviates from one for the effective radial components $P^{av}_{\ell}$, $\dot{P}^{av}_{\ell}$ and $P^{KH}_{l}$, $\dot{P}^{KH}_{l}$, described in Sec. II.1, because they are obtained from the Dirac equation. To illustrate this, in Table 2 we reproduce the values of the deviation factor
$F(l)=R_{MT}^{2}(\dot{P}_{l}(R_{MT})\,P^{\prime}_{l}(R_{MT})$
$-\dot{P}^{\prime}_{l}(R_{MT})\,P_{l}(R_{MT}))-1,$ (10)
for the avD and KH radial basis functions. For the nonrelativistic (Schrödinger) functions we have $F(l)\equiv 0$. In practice, as shown in Table 2 we find $F(l)\neq 0$. The deviations are the largest (F(p)=0.13F(p)=-0.13) for the 6pp radial functions in the avD-basis set. In the KH basis set $F(l)$ are smaller, reaching only the value $F(p)\approx 7\cdot 10^{-4}$ for $6p$-states of Ac. However, even such deviations can lead to a sizeable inaccuracy in determination of the equilibrium lattice constants and bulk moduli, which should be prevented.
It is not difficult to make corrections in LAPW method, accommodating it with the fact that $F(l)\neq 0$. The corrections involve the precise determination of the $a_{l}$ and $b_{l}$ coefficients and the matrix elements for the spherically symmetric component of the potential. In particular, Eq. (10b) and Eq. (10d) of Ref. 8 should be replaced with
$a^{n}_{l}=\frac{1}{\triangle}(j^{\prime}_{l}(k_{n}R_{MT})\,\dot{P}_{l}-j_{l}(k_{n}R_{MT})\,\dot{P}^{\prime}_{l}),$ (11a)
$b^{n}_{l}=\frac{1}{\triangle}(j_{l}(k_{n}R_{MT})\,P^{\prime}_{l}-j^{\prime}_{l}(k_{n}R_{MT})\,P_{l}),$ (11b)
where
$\triangle=R_{MT}^{2}(\dot{P}_{l}\,P^{\prime}_{l}-P_{l}\,\dot{P}^{\prime}_{l})\neq 1.$ (12)
The value for $\gamma^{l}$ given in Eq. (16b) of Ref. [8], should also be rewritten. In the notation of Ref. [8] the following symmetric form can be obtained
$\gamma^{l}$ $=$ $\frac{1}{2}\left\{a_{l}(\vec{k}_{n})b_{l}(\vec{k}_{m})+a_{l}(\vec{k}_{m})b_{l}(\vec{k}_{n})\right$. (13)
$+\left.\frac{1}{R^{2}_{MT}}(j^{\prime}_{l}(n)j_{l}(m)+j_{l}(n)j^{\prime}_{l}(m))\right\}.$
Here the second part with the Bessel functions comes from the MT-sphere boundary integration of the kinetic energy performed for the symmetrization of the expression for the matrix elements of kinetic energy, i.e. it appears due to the replacement of $\vec{k}_{n}\vec{k}_{n}U$ (or $\vec{k}_{m}\vec{k}_{m}U$) with $\vec{k}_{n}\vec{k}_{m}U$, Eq. (16a) of [8].
Table 2: Deviation of the factors $F(l)$ from zero, Eq. (10), for the avD, Eq. (3), and the canonical KH [22, 23] basis functions in fcc Th, fcc Ac, bcc Np, cubic ThO2 [Th(2), O(2)] and UO2 [O(3)], underlying the importance of the corrections for LAPW matrix elements, Eqs. (11a), (11b) and Eq. (13).
basis $s$ $p$ $d$ $f$
Th avD $1.2\cdot 10^{-4}$ 0.129-0.129 $4.0\cdot 10^{-4}$ $1.8\cdot 10^{-4}$
Th KH $1.1\cdot 10^{-4}$ $2.0\cdot 10^{-4}$ $7.9\cdot 10^{-5}$ $1.4\cdot 10^{-4}$
Th(2) avD $5.1\cdot 10^{-4}$ $3.3\cdot 10^{-3}$ $8.5\cdot 10^{-4}$ $2.4\cdot 10^{-4}$
Th(2) KH $5.1\cdot 10^{-4}$ $2.4\cdot 10^{-4}$ $2.1\cdot 10^{-4}$ $2.1\cdot 10^{-4}$
O(2) avD $6.7\cdot 10^{-5}$ $7.3\cdot 10^{-5}$ $1.1\cdot 10^{-5}$ $1e^{-6}$
O(2) KH $6.6\cdot 10^{-5}$ $7.1\cdot 10^{-5}$ $1.4\cdot 10^{-5}$ $1.3\cdot 10^{-5}$
U avD $5.2\cdot 10^{-4}$ $1.5\cdot 10^{-3}$ $8.3\cdot 10^{-4}$ $2.6\cdot 10^{-4}$
U KH $5.2\cdot 10^{-4}$ $2.6\cdot 10^{-4}$ $2.1\cdot 10^{-4}$ $2.9\cdot 10^{-4}$
O(3) avD $6.2\cdot 10^{-5}$ $7.8\cdot 10^{-5}$ $1.3\cdot 10^{-5}$ $2.0\cdot 10^{-6}$
O(3) KH $6.5\cdot 10^{-5}$ $7.7\cdot 10^{-5}$ $1.6\cdot 10^{-5}$ $1.4\cdot 10^{-5}$
Ac avD $1.0\cdot 10^{-4}$ 0.126-0.126 $3.6\cdot 10^{-4}$ $1.6\cdot 10^{-4}$
Ac KH $7.5\cdot 10^{-4}$ $6.8\cdot 10^{-4}$ $4.5\cdot 10^{-4}$ $3.9\cdot 10^{-4}$
Np avD $2.6\cdot 10^{-4}$ 0.039-0.039 $5.5\cdot 10^{-4}$ $2.9\cdot 10^{-4}$
Np KH $2.6\cdot 10^{-4}$ $2.4\cdot 10^{-4}$ $1.3\cdot 10^{-4}$ $3.1\cdot 10^{-4}$
II.3 Calculated spin-orbit coupling constants, special treatment for $6p$-states
In this section we consider how the new (avD) basis functions affect the values of the SO energy splittings. As a test exercise we first calculate SO coupling constants and energy splittings for relativistic atoms of Ac, Th, U and Np with the PBE variant of DFT for exchange and correlations [27].
In atomic units the SO coupling constant $\zeta(l)$, defined by the radial function $P_{l}(r)$, can be found as
$\zeta(l)=\frac{1}{2c^{2}}\int_{0}^{\infty}dr\,P_{l}^{2}(r)\,\frac{1}{r}\frac{dV}{dr},$ (14)
where $V(r)$ is the radial dependence of the Coulomb potential. The corresponding SO operator is
$H^{SO}=\zeta(l)\,\hat{L}\hat{S}$, (15)
with energy splitting
$\triangle_{SO}(l)=\zeta(l)\frac{2l+1}{2}.$ (16)
As $P_{l}$ we consider either $P_{\ell}^{av}$ (the avD basis) or $P_{l}^{KH}$ (the KH basis). Comparing $\triangle_{SO}(l)$, Eq. (16), with the actual splitting $\triangle E(l)$ of energy components in relativistic atom, we can conclude which basis set (avD or KH) gives a better description of the SO splitting. For the KH-basis, during the self-consistent procedure all core electron shells were obtained according to the fully relativistic Dirac approach whereas all valence electron shells according to the KH-equations [22, 23], as is done in LAPW calculations. The calculated values of the SO coupling for $6d$ states are listed in Table 3, for $5f$ states in Table 4 and for $6p$ states in Table 5. Since for the avD-basis set we calculate individual Dirac radial components, we also quote the individual SO couplings $\zeta(l)$ for them, i.e. for $d_{3/2}$, $d_{5/2}$ states in Table 3, for $f_{5/2}$, $f_{7/2}$ in Table 4 and for $p_{1/2}$, $p_{3/2}$ states in Table 5.
Comparing $\triangle_{SO}(l)$ with $\triangle E(l)$ in Table 3 for $d$-states and Table 4 for $f$-states shows that in all cases the calculated SO constants $\zeta^{avD}(\ell)$, based on $P_{\ell}^{av}$-functions, give much better energy differences $\triangle_{SO}^{avD}(d)$, $\triangle_{SO}^{avD}(f)$ than $\triangle_{SO}^{KH}(d)$, $\triangle_{SO}^{KH}(f)$, based on the KH functions $P_{l}^{KH}$. This is directly related to the behavior of the avD and KH radial functions close to the nuclear region, where the functions $P_{\ell=2}^{av}$ ($P_{\ell=3}^{av}$) are systematically larger than $P_{l=2}^{KH}$ ($P_{l=3}^{KH}$), Fig. 3. Although the difference between the $d$- and $f$- basis functions in the whole region is rather small, Table 1, the larger values of $P_{\ell=2}^{av}$ ($P_{\ell=3}^{av}$) is the decisive factor which finally leads to larger SO coupling constants and better values for the energy splittings.
Figure 3: Averaged radial functions $\bar{P}_{\ell}(r)$ (a) for $6d$ ($\ell=2$), and (b) for $5f$ ($\ell=3$) states and the corresponding canonical KH radial function $\bar{P}_{l}(r)$ [22, 23] close to the nuclear region.
The situation however is changed for the SO interactions of $6p$ states. The reason for this is a very different radial dependence of the $p_{1/2}$ and $p_{3/2}$ radial components in the nuclear region, shown in Fig. 1 and Fig. 2, and discussed earlier in Sec II.1. In the neighborhood of the nucleus, important for SOC, the $p_{1/2}$ component is very large (even singular for the point nucleus), and as a result, $\zeta_{1/2}$ calculated with the $p_{1/2}$ function is more than six times larger than $\zeta_{3/2}$ calculated with the $p_{3/2}$ component, Table 5. Even after the averaging between two $p$-components according to Eq. (3), the calculated SOC constant $\zeta^{avD}(p)$ overestimates approximately twice the actual SO-splitting, i.e. for the avD-basis $\triangle_{SO}(p)\approx 2\times\triangle E(p)$, where $\triangle E(p)=E(6p_{3/2})-E(6p_{1/2})$ is the actual splitting. The use of KH $6p$ radial functions improves the situation but it is also far from being ideal. Although in the nuclear region the KH-radial $6p$ functions $P_{l=1}^{KH}$ [22, 23] is appreciably smaller than the corresponding avD-function $P_{\ell=1}^{av}$, Fig. 2A, its SO coupling constant $\zeta^{KH}(p)$ remains large. Inspection of Table 5 shows that $\triangle_{SO}^{KH}(p)$ overestimates the actual values $\triangle E(p)$ by 17% for Ac, 19% for Th and 27% for Np. Earlier the problem of overestimated SO coupling effects was noticed e.g. in Ref. [6]. The situation is aggravated by large absolute values of the $6p$-splittings (6.7-9.5 eV), which have a large impact on the band structure calculations.
From the table 5 we can conclude that the actual energy splittings $\triangle E(p)$ between the $p_{1/2}$ and $p_{3/2}$ states are better approximated by the SO coupling constant $\zeta(p_{3/2})$ calculated using a single radial component $p_{3/2}$, for which $\triangle_{SO}^{*}(p)=3\zeta(p_{3/2})/2$. For example, for Ac we obtain $\triangle_{SO}^{*}(p)=6.09$ eV, for Th $\triangle_{SO}^{*}(p)=7.12$ eV etc. Although the values $\triangle_{SO}^{*}(p)$ are slightly smaller than the real energy differences $\triangle E(p)$, they approximate $\triangle E(p)$ better (i.e. 9119-11% vs 172717-27%) than the KH radial functions $P_{l=1}^{KH}$. Although in this Section we have considered the results with the PBE variant of DFT, the same conclusions can be drawn for other DFT functionals. Therefore, in the following for the LAPW calculations of SOC constants $\zeta(p)$ for the $6p$ semicore band states we will use only $p_{3/2}$ radial component. For the LAPW calculations of the SO couplings $\zeta(d)$, $\zeta(f)$ for the $d$ and $f$ valence states we use the standard averaged radial components, Eq. (3), because, as discussed above, they give good approximations of the energy splittings. Our results with such SO coupling constants are reproduced in Tables 69 below.
Table 3: Calculated SO coupling constants $\zeta$ (in eV) with the $6d_{3/2}$ and $6d_{5/2}$ radial functions in atoms. $\zeta(d)$ is the averaged value for two basis sets: avD-basis, Eq. (3), and KH-basis, [22, 23]. $\triangle_{SO}(d)$ (in eV) is the corresponding SO energy splitting, whereas $\triangle E=E(6d_{5/2})-E(6d_{3/2})$ (in eV) is the difference according to the fully relativistic Dirac atomic calculation.
basis $\zeta(d_{3/2})$ $\zeta(d_{5/2})$ $\zeta(d)$ $\triangle_{SO}(d)$ $\triangle E(d)$
Ac avD 0.182 0.142 0.158 0.394 0.372
Ac KH 0.125 0.313 0.372
Th avD 0.250 0.196 0.216 0.541 0.510
Th KH 0.182 0.456 0.510
U avD 0.292 0.224 0.250 0.624 0.587
U KH 0.212 0.531 0.587
Np avD 0.254 0.188 0.214 0.534 0.502
Np KH 0.230 0.574 0.502
Table 4: Calculated spin-orbit coupling constants $\zeta$ (in eV) with the $5f_{5/2}$ and $5f_{7/2}$ radial functions in atoms. $\zeta(f)$ is the averaged value for two basis sets: avD-basis, Eq. (3), and KH-basis, [22, 23]. $\triangle_{SO}(f)$ (in eV) is the corresponding SO energy splitting, whereas $\triangle E=E(5f_{5/2})-E(5f_{7/2})$ (in eV) is the difference according to the fully relativistic Dirac atomic calculation.
basis $\zeta(f_{5/2})$ $\zeta(f_{7/2})$ $\zeta(f)$ $\triangle_{SO}(f)$ $\triangle E(f)$
Th avD 0.198 0.180 0.188 0.658 0.650
Th KH 0.164 0.573 0.650
U avD 0.276 0.254 0.263 0.922 0.909
U KH 0.242 0.848 0.909
Np avD 0.292 0.264 0.276 0.967 0.953
Np KH 0.283 0.992 0.953
Table 5: Calculated SO coupling constants $\zeta$ (in eV) with the $6p_{1/2}$ and $6p_{3/2}$ radial atomic functions. $\zeta(p)$ is the averaged value for two basis sets: avD basis, Eq. (3), and KH-basis, [22, 23]. $\triangle_{SO}$ (in eV) is the corresponding SO energy splitting, $\triangle^{*}_{SO}=\triangle_{SO}(p_{3/2})$ (in eV) is the energy splitting for $\zeta(p_{3/2})$, whereas $\triangle E=E(6p_{3/2})-E(6p_{1/2})$ (in eV) is the actual difference according to the fully relativistic Dirac atomic calculation.
basis $\zeta(p_{1/2})$ $\zeta(p_{3/2})$ $\zeta(p)$ $\triangle^{*}_{SO}$ $\triangle_{SO}$ $\triangle E$
Ac avD 26.22 4.06 8.17 6.09 12.25 6.69
Ac KH 5.22 7.84 6.69
Th avD 32.16 4.75 9.78 7.12 14.66 7.80
Th KH 6.18 9.26 7.80
U avD 42.98 5.58 12.28 8.36 18.42 9.26
U KH 7.40 11.10 9.26
Np avD 47.40 5.66 13.05 8.49 19.57 9.53
Np KH 8.06 12.08 9.53
In addition to the SO couplings occurring inside the MT-sphere region, we have examined the SO effect in the interstitial region (IR). The matrix elements of the SOC there are given by
$\langle\phi_{p}|V^{SO}|\phi_{j}\rangle=\frac{i}{4c^{2}}\sum_{\vec{K}}F(\vec{K}_{j}-\vec{K}_{p}+\vec{K})\,V_{\vec{K}}$
$\left[\vec{K}\times(\vec{k}+\frac{1}{2}(\vec{K}_{j}+\vec{K}_{p}))\right]\vec{\sigma},$ (17)
where $\vec{K}_{j}$, $\vec{K}_{p}$ are the corresponding reciprocal lattice vectors, $V_{\vec{K}}$ is the Fourier component of the potential in IR, $\vec{\sigma}$ are the Pauli matrices, $F(\vec{K})$ are the standard LAPW integral of $\exp(i\vec{K}\vec{R})/v$ in IR. Our calculations indicate that the effect of additional SOC in the interstitial region, Eq. (17), is negligible. This is related to the fact that the variations of the total potential in IR are very small in comparison with the changes caused by nuclei. Therefore, the interstitial region can be safely considered as nonrelativistic.
II.4 Averaged account of small radial components
We have also estimated the influence of the small components $Q_{-\kappa}$ on the total electron density in the neighborhood of the nucleus (here $\kappa^{S}=-\kappa^{L}\equiv-\kappa$). In dealing with the $Q_{-\kappa}$ radial functions we should have in mind the following: the $Q_{-\kappa}$ functions are relatively large only in the neighborhood of the nuclear region, falling to a very small value of the order of $1/c$ close to the MT-sphere boundary (at $R_{MT}$) and in the interstitial region, Fig. 4. Since in the nuclear region the potential is spherically symmetric, we can make such an estimation, adding to the electron density $\rho(r)$ the spherically symmetric contribution $\rho_{Q}(r)$, associated with small components $Q_{-\ell}$ and $Q_{\ell+1}$, as described in more detail in Appendix B. Here the most important contribution comes from the fully occupied $6p_{1/2}$ and $6p_{3/2}$ states, accounting for more than 96% the additional $Q$-density. The results for the average valence electron density at finite nuclei ($\rho_{nuc}$) are given in Tables 69.
Figure 4: Radial function (avD) $P_{\ell=1}(r)$ of $6p$-states constructed from two large components ($P_{1}$ and $P_{-2}$, Eq. (3)) of the $6p_{1/2}$ and $6p_{3/2}$ states of thorium, and two associated small components $Q_{-1}$ and $Q_{2}$ (a) close to the nuclear region, and (b) on a larger scale.
Inspection of Tables 69 shows that the increase of the valence electron density is substantial, ranging from 2.3 to 4.3 times. However, the whole effect for the total electron density is diminished by a large contribution from the core electron states. The core states in actinides include in total 22 electron shells (from $1s$ to $5d$ and $6s$) accommodating 80 electrons with large values of electron density from the large and small components in the nuclear region. Their contribution, ranging from $4.3\cdot 10^{6}$ for Ac to $5.4\cdot 10^{6}$ for U (in atomic units, au-3), is approximately four orders of magnitude larger than the valence density. As a result, the overall relative change of the electron density is less than $10^{-4}$. Nevertheless, as quoted in Tables 69, there is a small influence on equilibrium parameters which in some cases can shift the equilibrium lattice constant by 0.007 Å and the bulk modulus by 8 GPa, Sect. III. The effect possibly should be also taken into account for the calculation of electron characteristics related to the nuclear spectroscopies, such as the nuclear magnetic resonance (NMR), Mössbauer spectroscopy etc.
III Application to actinides
For calculation of the exchange-correlation potential and the exchange-correlation energy contribution within the DFT approach, we have used (1) the Perdew-Burke-Ernzerhof (PBE) scheme [27] of the generalized-gradient approximation (GGA), (2) PBEsol [30] which is a variant of PBE, and (3) the local density approximation (LDA) with the standard ($V_{exc}\sim-\rho^{1/3}$) exchange [31] and the PW-correlation [32]. For the band structure calculations we have used the Moscow-FLAPW method [11, 28] which has been widely used by us before for the study of chemical bonding elemental solids and compounds.
The technical parameters of numerical calculations were as follows. For fcc Ac and fcc Th, in most cases the number of augmented plane waves was 137 and 274 (with SO), with $R_{MT}K_{j}\sim 10$, where $\vec{K}_{j}$ is the maximal wave vector, and $R_{MT}$ is the radius of the MT-sphere. For CaF2 cubic structures, the basis sets were 307 and 614 (with SO) for ThO2 ($R_{MT}K_{j}\sim 8.7$), 387 and 774 (with SO) for UO2 ($R_{MT}K_{j}\sim 9.5$). For the KH (canonical) basis two sets of $6p$ local functions ($u$ and $\dot{u}$) as described in Ref. [11]. Therefore, for the KH-calculations the total number of basis functions is 143 and 286 (with SO) for Ac and Th, 313 and 626 (with SO) for ThO2, and 393 and 786 for UO2. The following radii of MT-spheres have been used: for Ac 3.46 a.u. for all DFT functionals (i.e. LDA, PBEsol, PBE); for Th 3.05 a.u. (LDA), 3.12 a.u. (PBEsol and PBE); for ThO2 2.25 a.u. for both Th and O for all DFT functionals; for UO2 2.138 a.u. (LDA), 2.168 a.u. (PBEsol, PBE) for both U and O. It is worth noticing that within the chosen DFT functional we used the same $R_{MT}$ for all variants of the different radial functions (i.e. avD, avD(Q), KH with and without SOC). The dependence of the results of the new treatment on the chosen $R_{MT}$ is briefly discussed in Appendix C. The maximal number of $k$-points in the irreducible part (IP) of the Brillouin zone (BZ) for elemental actinides was 1505 ($\sim 70000$ for the whole BZ). For ThO2 and UO2 we have used a set of 240 $k$-points in IP of BZ ($\sim 11500$ for the whole BZ). The maximal value of the LAPW plane-wave expansion and the non-spherical density decomposition was $L_{max}=8$. We have taken into account the finite size of nuclei and used the tetrahedron method for the linear interpolation of energy between $k$-points [29]. The number of radial points inside the MT sphere region was increased to 4000-4200 for actinides and 900-1000 for oxygen. The enlarged number of radial points is analogous to the increase of the quality of the basis set, and this is certainly required for the actinides with 80 core electrons, which is a very large quantity. The calculations with the KH basis sets have been performed with additional localized atomic $p$-basis functions as described in Ref. [11]. New radial basis functions $P_{\ell}^{av}$ (avD) have turned out to work well even without them, and all calculations marked in the Tables as avD or avD(Q) (with and without SO) have been carried out without localized functions.
We employed the full treatment of the SO effects, which effectively doubles the dimension of the basis set. This option makes difference with other FLAPW calculations of actinides [6, 7] where the SO coupling was incorporated at the second variational level [24, 25], which introduces certain uncontrolled approximations [10].
The results of our calculations are listed in Table 6 for fcc Ac, Table 7 for fcc Th, Table 8 for cubic ThO2, and Table 9 for cubic UO2. ThO2 and UO2 are crystallized in the CaF2 structure. As discussed in the Introduction, most calculations of actinides [2, 4] and their oxides [5, 3, 33, 34] have been performed with the full potential LMTO method. Although nowadays the full potential LAPW (FLAPW) study of actinides [6, 7] and dioxides [35, 36] are also available, the new feature of the present FLAPW calculations is the complete treatment of the SO couplings. It is also worth mentioning that there are several theoretical studies of UO2 with correlation effects (Hubbard repulsion) [37, 38, 39, 40, 41, 42, 43]. Such an approach however lies beyond the scope of the present work, which focuses on the peculiarities of the inclusion of relativistic effects.
All calculations in the present work are performed for two various basis sets (with the canonical KH radial functions and averaged Dirac [avD] functions, Sec. II.1). For the avD and avD(Q) basis sets we have used the corrected values for $a_{l}^{n}$, $b_{l}^{n}$, Eq. (11a), Eq. (11b), and Eq. (13) for $\gamma^{l}$ for the matrix elements in the spherical ($L=0$) component. For the KH basis set we have adopted the values given in Eq. (16a) and Eq. (16b) of Ref. [8]. In addition, we have carried out calculations with and without the SO coupling. For the calculation of the SO coupling constant of the $6p$ semicore states in the avD or avD(Q) bases we used the $6p_{3/2}$ large component as described in Sec. II.3, and for the KH basis the canonical averaged radial $6p$ component, which overestimates the SO energy splitting, see more details in Sec. II.3. For the SO coupling constants of other valence states (i.e., $d$-, $f$-, and higher $\ell$- states), we employed averaged radial functions, which, however, as discussed in Section II.3, differ somewhat in the avD and KH schemes. Finally, for the avD basis set we have tried two variants – with and without additional contribution to the electron density from small ($Q$) components of Dirac solutions, Sec. II.4, marked below as avD and avD(Q).
Inspection of Tables 69 shows that even within the same DFT functional (LDA, PBE or PBEsol), various inclusions of relativistic effects lead to very different results for the equilibrium lattice constants and bulk modulii. In particular, the largest variation of $a$ reaches 0.147 Å for fcc Ac (LDA) although in the case of ThO2 it is only 0.019 Å for LDA and 0.01 Å for PBE and PBEsol. The largest difference in $B$ reaches 26.2 GPa for fcc Th (LDA) and 24 GPa for UO2 (LDA), although, for example, for Ac (PBE) it is only 2.4 GPa. Inclusion of the SO coupling leads to smaller lattice constants for fcc Ac and Th, but to larger ones for UO2. As a rule, the SO coupling results in larger bulk moduli, but in some cases they practically do not change (Ac, PBE and PBEsol; ThO2, all DFT) or even get smaller (Ac, LDA, PBE, PBEsol with avD/avD(Q); or UO2 with all DFT variants).
The opposing trends are also found for the avD/avD(Q) and KH basis sets. In some cases the use of the KH functions leads to smaller lattice constants (fcc Ac and Th), but in other instances (ThO2 or UO2) it gives larger values of $a$. The bulk moduli calculated with the KH functions can be larger (Th, all DFT variants; ThO2, LDA), but also smaller than $B$ found with the avD/avD(Q) variants (UO2, all DFT functionals). The other characteristics of the band structure are also susceptible to different treatment of relativistic effects. For example, the gap $E_{g}$ of forbidden states in ThO2 changes by 0.25 eV ($\sim 5$%), Table 8.
Due to the inclusion of the density of the small component ($Q$) the valence electron density at the nucleus is substantially increased: 2.4–2.5 times for Ac and Th, 3.9 for Th in ThO2 and 4.1–4.3 for U in UO2. However, because of the very large contribution of the core electrons the total variation of electron density at nucleus appears to be small. The relative change is only $4\cdot 10^{-5}$ for Ac and Th in ThO2, and $\sim 4.6\cdot 10^{-5}$ for elemental Th and U in UO2. Generally, the inclusion of $Q$-components gives little difference in results. However, in some exceptional cases it accounts for 0.007 Å change in $a$ (fcc Ac, LDA with SO), and 8 GPa in $B$ (ThO2, LDA).
Table 6: Results of LAPW calculations for fcc structure of elemental actinium (Ac) with the averaged Dirac (avD) and Koelling-Harmon [22] (KH) radial basis functions, with SO (marked by *) and without it. avD(Q) stands for calculations with small components for electron density; $a$ is the equilibrium lattice constant (in Å), $B$ is the bulk modulus (in GPa), $\rho_{nuc}^{Ac}$ is the valence electron density inside the nucleus (in a.u.-3). Experimental data: $a=5.315$ Å [44], estimated $B=24.5$ GPa [46].
DFT Basis SO $a$ (Å) $B$, GPa $\rho_{nuc}^{Ac}$
LDA avD 5.576 28.2 130.7
LDA avD(Q) 5.573 27.9 316.4
LDA avD * 5.540 27.6 131.0
LDA avD(Q) * 5.533 26.4 316.2
LDA KH 5.496 31.5 128.1
LDA KH * 5.429 27.2 127.9
PBE avD 5.756 24.6 122.9
PBE avD(Q) 5.754 24.1 306.1
PBE avD * 5.723 24.1 124.0
PBE avD(Q) * 5.718 23.5 306.1
PBE KH 5.682 24.6 120.3
PBE KH * 5.611 25.9 121.1
PBEsol avD 5.633 25.9 126.9
PBEsol avD(Q) 5.629 25.0 311.5
PBEsol avD * 5.592 25.4 128.0
PBEsol avD(Q) * 5.586 24.9 312.0
PBEsol KH 5.553 26.0 125.1
PBEsol KH * 5.479 27.9 125.8
Table 7: Results of LAPW calculations for fcc structure of elemental thorium (Th) with the averaged Dirac (avD) and Koelling-Harmon [22] (KH) radial basis functions, with SO (marked by *) and without it. avD(Q) stands for calculations with small components for electron density; $a$ is the equilibrium lattice constant (in Å), $B$ is the bulk modulus (in GPa), $\rho_{nuc}^{Th}$ is the valence electron density inside the nucleus (in a.u.-3). Experimental data: $a=5.0845$ Å, $B=58$ GPa [45].
DFT Basis SO $a$ (Å) $B$, GPa $\rho_{nuc}^{Th}$
LDA avD 5.015 74.1 165.2
LDA avD(Q) 5.013 74.4 394.8
LDA avD * 4.996 63.7 165.5
LDA avD(Q) * 4.992 61.5 394.0
LDA KH 4.956 82.7 162.2
LDA KH * 4.910 87.7 161.6
PBE avD 5.142 57.0 156.9
PBE avD(Q) 5.140 56.2 384.7
PBE avD * 5.119 55.6 157.5
PBE avD(Q) * 5.115 54.8 385.2
PBE KH 5.066 57.7 154.2
PBE KH * 5.009 63.0 155.0
PBEsol avD 5.054 58.6 162.9
PBEsol avD(Q) 5.049 58.8 390.9
PBEsol avD * 5.026 60.1 163.2
PBEsol avD(Q) * 5.021 59.1 391.2
PBEsol KH 4.971 61.7 160.0
PBEsol KH * 4.921 69.2 159.6
Table 8: Results of LAPW calculations of uranium dioxide ThO2 (CaF2 structure) with with the averaged Dirac (avD) and Koelling-Harmon [22] (KH) radial basis functions, with SO (marked by *) and without it. avD(Q) stands for calculations with small components for electron density; $a$ is the equilibrium lattice constant (in Å), $B$ is the bulk modulus (in GPa), $\rho_{nuc}^{Th}$ is the valence electron density of Th inside the nucleus (in a.u.-3). Experimental data [47]: $a=5.6001$ Å, $B=198$ GPa.
DFT Basis SO $a$ (Å) $B$, GPa $E_{g}$, eV $\rho_{nuc}^{Th}$
LDA avD 5.587 201.0 4.65 70.3
LDA avD(Q) 5.584 209.0 4.64 272.0
LDA avD * 5.592 208.6 4.54 70.3
LDA avD(Q) * 5.590 205.0 4.54 271.2
LDA KH 5.596 232.3 4.50 66.1
LDA KH * 5.603 228.5 4.38 65.9
PBE avD 5.686 198.5 4.69 65.8
PBE avD(Q) 5.686 198.6 4.69 265.2
PBE avD * 5.687 200.0 4.59 65.9
PBE avD(Q) * 5.687 200.0 4.59 264.4
PBE KH 5.692 198.4 4.53 61.7
PBE KH * 5.697 195.8 4.44 61.6
PBEsol avD 5.621 215.3 4.63 68.4
PBEsol avD(Q) 5.620 215.3 4.63 268.4
PBEsol avD * 5.622 216.7 4.53 68.4
PBEsol avD(Q) * 5.622 216.6 4.53 267.6
PBEsol KH 5.627 215.4 4.50 64.9
PBEsol KH * 5.632 211.8 4.39 64.8
Table 9: Results of LAPW calculations of uranium dioxide UO2 (CaF2 structure) with with the averaged Dirac (avD) and Koelling-Harmon [22] (KH) radial basis functions, with SO (marked by *) and without it. avD(Q) stands for calculations with small components for electron density; $a$ is the equilibrium lattice constant (in Å), $B$ is the bulk modulus (in GPa), $\rho_{nuc}^{U}$ is the valence electron density of U inside the nucleus (in a.u.-3). Experimental data [47]: $a=5.4731$ Å, $B=207$ GPa.
DFT Basis SO $a$ (Å) $B$, GPa $\rho_{nuc}^{U}$
LDA avD 5.317 279.7 84.5
LDA avD(Q) 5.316 279.2 361.2
LDA avD * 5.346 264.5 82.3
LDA avD(Q) * 5.345 263.8 355.8
LDA KH 5.332 275.6 78.0
LDA KH * 5.358 255.7 76.8
PBE avD 5.435 234.8 85.2
PBE avD(Q) 5.434 234.7 361.5
PBE avD * 5.468 224.3 83.0
PBE avD(Q) * 5.467 224.0 356.1
PBE KH 5.451 231.0 78.3
PBE KH * 5.481 216.8 77.6
PBEsol avD 5.365 259.8 89.9
PBEsol avD(Q) 5.364 259.2 368.0
PBEsol avD * 5.396 245.1 87.4
PBEsol avD(Q) * 5.395 244.8 362.2
PBEsol KH 5.381 255.4 83.0
PBEsol KH * 5.408 237.3 82.1
It is also worth mentioning that in contrast to Th, our DFT calculations of the fcc structure of Ac appreciably overestimate its lattice constant even for LDA. This however was also noticed e.g. in Ref. [48] ($a=5.503$ Å in LDA) and therefore, the effect should be attributed to the peculiarity of the band structure of this element. Our calculations indicate that this feature becomes more pronounced with increasing quality of the basis set. For example, decreasing the non-spherical components of electron density to $L_{max}=6$ in the avD(Q) basis leads to a smaller lattice constants: 5.536 Å ($\delta a=-0.037$ Å) in LDA, 5.736 Å (0.018-0.018 Å) in PBE, 5.597 Å (0.032-0.032 Å) in PBEsol. Further, decrease of the basis set to only 65 functions results in 5.500 Å (total $\delta a=-0.073$ Å in LDA) 5.608 Å (0.146-0.146 Å) in PBE, 5.539 Å (0.09-0.09 Å) in PBEsol, compare with Table 6. Possibly, some properties of the phonon spectrum and mean square displacements of atoms in solid Ac make the description using poor basis sets more adequate to the experimental data. Owing to its scarcity and radioactivity, the experimental bulk modulus of Ac is unknown. To the best of our knowledge in the literature there is only an estimated (not directly measured) value of 24.5 GPa, listed in Ref. [46]. There is also a theoretical estimation, $B=25.9$ GPa, performed on the basis of a tight-binding analysis [48]. Both values are in good correspondence with our PBE and PBEsol data, Table 6, which represent ab initio theoretical data for bulk moduli of Ac.
Finally, we would like to comment on our calculations of UO2, Table 9 and Fig. 5. In particular, it is often stated that in contrast to the experimental observations the plain band structure analysis predicts the metal character of this compound. This is not completely correct if the SO coupling is fully taken into account. As shown in Fig. 5, when the SO coupling is included, there is a small gap of 0.20.40.2-0.4 eV between the highest occupied and the lowest unoccupied $5f$ bands at the Fermi energy. The appearance of the gap in the $5f$ band spectrum can be understood as follows. In the U atom the $5f_{5/2}$ and $5f_{7/2}$ electron states are split by approximately 1 eV because of the SO interaction. In the UO2 compound each U atom is surrounded by 8 oxygen neighbors, and the corresponding crystal electric field (CEF) causes additional energy splittings according to the schemes [49],
$D_{J=5/2}\rightarrow\Gamma_{7}+\Gamma_{8},$
$D_{J=7/2}\rightarrow\Gamma_{6}+\Gamma_{7}+\Gamma_{8},$
In particular, the lowest $J=5/2$ level is split into a doublet ($\Gamma_{7}$) and a quartet ($\Gamma_{8}$). (The $5f$ CEF splittings can be traced at the $\Gamma$ point in Fig. 5.) A small overlap of $5f$ states of U provides the electron band structure, with the $\Gamma_{8}$ quartet giving rise to two lowest unoccupied $5f$ bands, but, as follows from Fig. 5, a small energy difference between the $5f$ states, originating from the split $\Gamma_{7}$ and $\Gamma_{8}$ states, is preserved. The gap is not clearly observed because the neighboring occupied and unoccupied $5f$ bands slightly overlap. However, the same effect was found in calculations of crystalline germanium [50], which is nevertheless considered a dielectric material.
Figure 5: The upper panel of the calculated band structure of UO2 with the spin-orbit coupling, PBE calculation. The highest occupied $5f$ electron band is shown by solid line, lowest unoccupied $5f$ bands by dashed lines. The vertical gap with $\triangle E$ from 0.20.40.2-0.4 eV is visible.
IV Conclusions
Within the LAPW method we have presented a few ways to include the relativistic effects more completely and consistently: (1) we have used new radial basis functions $P_{\ell}^{av}$ and $\dot{P}_{\ell}^{av}$, which are obtained by finding the large components $P_{\kappa=\ell}$, $P_{\kappa=-\ell-1}$ of the Dirac solutions independently for the $j=\ell-1/2$ and $j=\ell+1/2$ states and then averaging them explicitly by means of Eq. (3). We have found that new radial basis functions $P_{\ell}^{av}$ bring more stability to the self-consistent-field procedure, performing quite well even without additional $6p$ semicore basis functions used e.g. in Ref. [11], but also without the $6p_{1/2}$ function used in the second variation SO treatment of Refs. [24, 25]; (2) we have corrected the LAPW expressions for $a_{l}^{n}$, $b_{l}^{n}$, Eq. (11a), Eq. (11b), and for the matrix elements in the spherical ($L=0$) component of the total potential, using Eq. (13) for $\gamma^{l}$. The canonical expression for $\gamma^{l}$ given in Eq. (16a) and Eq. (16b) of Ref. 8 implicitly uses Eq. (9), which is valid only for the non-relativistic radial solutions $u_{l}$ and $\dot{u}_{l}$, Sec. II.2; (3) for the calculation of the SOC constants for the semicore $6p$-states we have used the $P_{\kappa=-2}$ large component of the Dirac solution for the $6p_{3/2}$ states, which, as demonstrated in Sec. II.3, gives better approximation for the actual energy splittings. For $6d$ and $5f$ levels, the SOC constants are calculated with the averaged components $P_{\ell}^{av}$, which describe the SO energy splittings of these states adequately; (4) we have taken into account the additional electron density ($Q$-density) close to the nuclear region by including the contributions from the small components of the $6p_{1/2}$ and $6p_{3/2}$ and other states.
Based on our calculations, the discussed difference in the treatment of relativistic effects can result in uncertainties up to 0.15 Å for lattice constants and to 26 GPa for bulk moduli even within the same chosen DFT functional (LDA, PBE or PBEsol), Tables 69.
Due to the inclusion of the density from the small component ($Q$) the valence electron density at the nucleus is substantially increased (2.3–4.3 times). However, the final effect for electron density is not well pronounced because of a large background contribution from the core electrons, with the relative change of the order of $10^{-4}$. Nevertheless, in some instances the inclusion of the small component ($Q$) accounts for 0.007 Å change in lattice constants and 8 GPa in bulk modulus. We have also examined the SO coupling in the interstitial region, using Eq. (17) for the matrix elements, but the effect appears to be negligible. Unfortunately, as discussed in Sec. III it is not possible to conclude on the direction of the changes (i.e. increase or decrease of $a$ or $B$) when different relativistic treatments are involved: in different materials the trends are opposite.
Our calculations include the full scale account of the SO coupling for valence states, which differs from other studies where the SO coupling is omitted or treated as a second variation step with certain approximations [24, 25, 10]. The calculated bulk moduli for Ac are in good correspondence with the experimental estimation of $B=24.5$ GPa ($0.25\times 10^{6}$ kg/cm2) made in Ref. [46] and the only theoretical value $B=25.9$ GPa, obtained on the basis of a tight-binding analysis in Ref. [48]. As for the calculation of the band structure of UO2, it is worth noting that in the literature it is often wrongly considered as having metallic band structure. Our calculations of UO2 with the SO coupling clearly demonstrate that it has a small gap ($\sim 0.2-0.4$ eV) at the Fermi energy, Fig. 5. Although the upper $5f$ occupied band and the next valence $5f$ band in UO2 slightly overlap, the situation is completely analogous to the calculated band structure of germanium, [50], which is a well known dielectric.
Acknowledgements.
This research was supported by a grant of the Russian Science Foundation (Project No 24-12-00053).
Appendix A
In general, the factorized 4-spinor of the solution of the Dirac equation in the central field is written as
$\psi(r,\Omega)=\frac{1}{r}\left(\begin{array}[]{c}P_{\kappa}(r)\,\xi_{\kappa,m}(\Omega)\\ Q_{-\kappa}(r)\,\xi_{-\kappa,m}(\Omega)\end{array}\right),$ (21)
where $\Omega=(\Theta,\phi)$ stands for two polar angles, and $\xi_{\pm\kappa,m}(\Omega)$ are the angular two-spinors, referring to the large ($L$) and small ($S$) component [21]. Note that $\kappa^{L}=-\kappa^{S}$. For the $p_{1/2}$-states $\kappa^{L}=1$ and the angular dependence of the large component spinor $\xi^{L}$ is described by $\ell=1$ angular functions [i.e spherical harmonics $Y_{\ell=1}^{m}(\Omega)$], as expected for $p$-states. For the small component however we have $\kappa^{S}=-\kappa^{L}=-1$, implying that the angular dependence of $\xi^{S}$ is constructed with $\ell^{S}=0$ functions (i.e. the $s$-function $Y_{\ell=0}^{m=0}=1/\sqrt{4\pi}$). For the $p_{3/2}$-states $\kappa^{L}=-2$, and the angular dependence $\xi^{L}$ of the large component is given again by the $\ell=1$ spherical harmonics $Y_{\ell=1}^{m}(\Omega)$. However, the angular dependence of its small component $\xi^{S}$ is described by the $\ell^{S}=2$ coordinate functions $Y_{\ell=2}^{m}(\Omega)$ (i.e. by $d$-functions) [21].
Appendix B
In the spherically symmetric potential the $j=\ell-1/2$ and $j=\ell+1/2$ levels are degenerate. If both levels are completely filled, which holds e.g. for the fully occupied $6p_{1/2}$ and $6p_{3/2}$ semicore states of the actinides, then, according to the generalized Unsöld theorem, the sum of squares of all angular components of the $j=\ell-1/2$ and $j=\ell+1/2$ state is an invariant independent of the polar angles. In this case the partial weights of the $j=\ell-1/2$ and $j=\ell+1/2$ levels in the total electron density are $w_{\ell,1}=\ell/(2\ell+1)$ and $w_{\ell,2}=(\ell+1)/(2\ell+1)$, correspondingly. Therefore, these weight factors can be used to correct the electron density, constructed from the averaged large component, by adding the extra terms associated with the small components $Q_{-\ell}(r)$ and $Q_{\ell+1}$ of these states, i.e.
$\rho_{\ell}(r)=\rho_{\ell,0}\left[P_{\ell}^{2}(r)+w_{\ell,1}Q_{-\ell}^{2}(r)+w_{\ell,2}Q_{\ell+1}^{2}(r)\right],$ (22)
where $\rho_{\ell,0}$ is the $\ell$-partial charge obtained for the averaged large radial component $P_{\ell}$. In LAPW method there are two radial functions – $P_{\ell}$ and $\dot{P}_{\ell}=dP_{\ell}/dE$. Therefore, in addition to $\rho_{\ell}(r)$, Eq. (22) we also take into account the electron density $\rho^{E}_{\ell}(r)$, associated with the energy derivatives $\dot{Q}_{-\ell}(r)$ and $\dot{Q}_{{\ell}+1}$ of the small components $Q_{-\ell}(r)$ and $Q_{{\ell}+1}$, i.e.
$\rho_{\ell,\,tot}(r)=\rho_{\ell}(r)+\rho^{E}_{\ell}(r),$ (23)
where
$\rho^{E}_{\ell}(r)=\rho^{E}_{\ell,0}\left[\dot{P}_{\ell}^{2}(r)+w_{\ell,1}\dot{Q}_{-\ell}^{2}(r)+w_{\ell,2}\dot{Q}_{\ell+1}^{2}(r)\right].\quad$ (24)
Here $\rho^{E}_{\ell,0}$ is the partial charge for $\dot{P}_{\ell}$. The electron densities (23) are involved in the calculation variants denoted as avD(Q) in tables 69. The total valence values of $\rho(r)$, averaged over the actinide nuclei of finite size, are denoted in the last column of these Tables as $\rho_{nuc}$. They can be compared with the electron densities obtained only with the large components $P_{\ell}$ and $\dot{P}_{\ell}$. Practically all increase (96%) of $\rho_{nuc}$ is due to the small components of the semicore $6p$ states.
Appendix C
We have studied the dependence of the results of new treatment on the chosen MT-radius $R_{MT}$ for fcc lattice of elemental thorium. The main peculiarity here is that we do not use the local $6p$ functions et al. As an example, we consider the equilibrium lattice constant $a$, given in Table 10.
Table 10: Dependence of the calculated equilibrium lattice constant value $a$ (Å) with new radial basis functions on chosen MT-sphere radius $R_{MT}$ for fcc thorium without SO coupling (the $Q$ variant includes small component valence density) and PBE density functional, with 137 LAPW basis functions. $R_{MT,1}=3.05$ a.u., $R_{MT,2}=3.12$ a.u., $R_{MT,1}=3.19$ a.u.
Basis $a$ (Å), $R_{MT,1}$ $a$ (Å), $R_{MT,2}$ $a$ (Å), $R_{MT,3}$
avD 5.138 5.142 5.144
avD(Q) 5.137 5.140 5.142
The calculated lattice constants $a$ demonstrate a weak dependence on $R_{MT}$. Indeed, the change $\triangle R_{MT}$ in going from $R_{MT,1}$ to $R_{MT,2}$, or from $R_{MT,2}$ to $R_{MT,3}$, which is 2.3% of $R_{MT}$, produces deviations of $a$ from $2\cdot 10^{-3}$ to $4\cdot 10^{-3}$ Å, which corresponds to a relative change of $\triangle a$ of $\sim 0.06$%, i.e., $\sim 40$ times less than $\triangle R_{MT}$. Such a change in $a$ is comparable to using different fitting schemes to find $a$ and $B$.
In our study we tend to use relatively large values of $R_{MT}$, which is related to our choice to work with a large number of radial points inside MT-sphere (4000-4200 points for actinides). The large number of radial points improves the accuracy of LAPW basis functions avoiding possible linear dependence. On the other hand, the large number of radial points improves the description of 80 core electrons for actinides, which is quite a big quantity requiring very accurate radial treatment.